Monochromatic 545-nm light is incident on a 15-μm-wide slit. If the diffraction pattern is cast on a screen 510 mm from the slit, what is the linear distance from the center of the pattern to the n = 4 dark fringe?
Answer #1

Wavel ength, \(\lambda=545 \times 10^{-9} \mathrm{~m}\), width of the slit, \(d=15 \times 10^{-6} \mathrm{~m}\)

\(n=4\), distance of the screen, \(D=510 \times 10^{-3} \mathrm{~m}\)

Line ar distance, \(l=\frac{n D \lambda}{d}\)

$$ \begin{array}{l} =\frac{4 \times 545 \times 10^{-9} \times 510 \times 10^{-3}}{15 \times 10^{-6}} \\ =0.074 \mathrm{~m} \end{array} $$